OpenStudy (anonymous):

If ln x + ln(x-4)= ln4x, then x= ?? The answer is 8...but can anyone help and show the working out?

OpenStudy (anonymous):

lna+lnb=ln(ab)

OpenStudy (anonymous):

lna=lnb => a=b

OpenStudy (lgbasallote):

Product Law... ln x(x-4) = ln4x e^ln4x = x(x-4) 4x = x(x-4) divide both sides by x.... 4 = x - 4 x = 8